OpenStudy (anonymous):

Please HELP me ): , i am stuck .. Suppose that y varies inversely with x, and y = 2 when x = 6. What is an equation for the inverse variation?

OpenStudy (dumbcow):

\[ y = \frac{k}{x}\] solve for k using given info \[2 = \frac{k}{6}\]

OpenStudy (anonymous):

/: im confused :(

OpenStudy (dumbcow):

the equation for inverse variation is: \[y = \frac{k}{x}\] where "k" is some constant so solve for k and you are done

OpenStudy (anonymous):

how do i solve for k ?

OpenStudy (dumbcow):

plug in x=6 and y=2

OpenStudy (anonymous):

@Luis_Rivera , help me .

OpenStudy (dumbcow):

am i and @Callisto not helping you ?

OpenStudy (anonymous):

your confusing me. im trying to plug it in, but then idk how to solve it. ): im just now learning how to do these problems.

OpenStudy (anonymous):

@adl_thoo , Show an attempt first

OpenStudy (anonymous):

I have :( im just learning how to do these. ive plugged it in, now how do i solve . thts where im stuck at.

OpenStudy (anonymous):

what did u "plug in " ?

OpenStudy (anonymous):

okay, i plugged in what they said & i got : 2=6/x

OpenStudy (anonymous):

look

OpenStudy (anonymous):

oka

OpenStudy (anonymous):

k

OpenStudy (anonymous):

o.O huh

OpenStudy (anonymous):

lol, that ,means ok

OpenStudy (anonymous):

lmao, i know it.

OpenStudy (anonymous):

@eSpeX , help me here ?

OpenStudy (espex):

What is it that you're not getting?

OpenStudy (anonymous):

how to solve the equation . im very confused .

OpenStudy (anonymous):

x=y/4 y=12/x y=4x y=x/12

OpenStudy (anonymous):

those are my answers, i cant seem to get any of them

OpenStudy (espex):

Okay, the problem tells you that the relation between x and y is an inverse relationship. This means that as y increases, x decreases.

OpenStudy (anonymous):

Right . soo ?

OpenStudy (espex):

As @dumbcow pointed out, your relationship is written as \[y=\frac{k}{x}\] and they give you the values of x and y.

OpenStudy (anonymous):

which means : 2=6/x

OpenStudy (espex):

So looking at your values and the potential answers, what (k) can you divide by x (6) that will give you y (2)?

OpenStudy (anonymous):

12 ?

OpenStudy (espex):

\[\boxed{\color{#E32}{\text{great job!}}}\]

OpenStudy (anonymous):

so my answers is B ?

OpenStudy (espex):

So looking at your answers, do you have something that looks like \[y=\frac{12}{x}\] ?

OpenStudy (espex):

Yes, nice work.

OpenStudy (anonymous):

yes, B :D

OpenStudy (anonymous):

thank you so much !

OpenStudy (espex):

You're welcome.